# Low-Loss Polarization-Maintaining Optical Router for Photonic Quantum Information Processing

### 12 Jan 2024 v1

@@ 1 Introduction
In this work, we propose and experimentally demonstrate a low-loss, polarization-maintaining EO router compatible with single photons. Our scheme is based on the polarization-maintaining operation of all optical components including custom birefringence-compensated EO crystals. This enables the construction of a router with a Mach-Zehnder interferometer (MZI) configuration without polarization compensators. Our experiment with classical laser pulses demonstrates 2-4% loss, 20 dB switching extinction ratio (SER), and 99% process fidelity in the polarization degree of freedom. The measured routing rise time is as fast as a standard bulk optic EOM (2.9 ns). We expect that our demonstration is an important step toward advanced polarization-encoded photonic quantum information technologies that rely on photon-polarization qubits and their entanglement.
@@ 2 Methods
where denotes the wavelength of the incident light; and are the length and width of the RTP crystals; () and () are the intrinsic refractive indices and EO coefficient for crystallographic () direction, respectively; is the voltage applied to the crystal. For identical applied fields to the two crystals, i.e., , one can achieve the same phase shift for orthogonal polarization states. Thus, the static and EO birefringence of the two crystals is compensated for each other, enabling a polarization-maintaining phase shift according to the applied electric field. We note that our scheme is generally applicable to a pair of identical EO crystals that is often fabricated for off-the-shelf thermally compensated Pockels cells. The other optical components of the router also preserve polarization states of light: in our squished configuration of the MZI, all optics components are operated with a nearly normal angle of incidence (AOI, ) to mitigate the polarization dependence (on S- and P-polarizations) in transmissivity, reflectivity, and phase shift[18, 19]. The interferometer arm without EOM is 16 mm longer than the other to compensate for the intrinsic group delay (relative to air) given by the EOM ( mm).
@@ 3 Results and Discussion
We constructed a polarization-maintaining router based on the scheme described above, using optical components with anti-reflection/high-reflection coatings optimized at the wavelength of nm. In this proof-of-concept experiment, we characterized our router via classical photodetection measurements of optical pulses from a diode laser at 785 nm; the performance on single-photon routing can be predicted by the classical measurements since our router utilizes the EO effect which is essentially independent of input optical power as shown in Eq. (1, 2) and produces no noise photons, as has been demonstrated in many quantum optics experiments [6, 20, 21, 22, 23]. A set of a polarizer (POL), a half-wave plate (HWP), and a quarter-wave plate (QWP) is used to prepare an arbitrary input polarization state. The router is operated at a 10 kHz repetition rate, which is limited by our high-voltage EOM driver; we expect that a 1 MHz repetition rate is available with our EOM that is compatible with off-the-shelf high-voltage drivers designed for standard Pockels Cells (used for polarization switching)[6, 16]. Our MZI is stable for several minutes which is limited by environmental disturbance. Nevertheless, the stability can be improved by miniaturizing the setup and/or utilizing matured active phase stabilization techniques to be applied for advanced quantum optics experiments.
@@
We measured optical power at the input port () and the output ports 1 and 2 ( and ) for different input polarization states and the applied voltage () to the EOM. Figure 2 shows the normalized output optical power ( and ) versus . The output power for horizontal (), diagonal (), and right-circular () polarized light changes according to the applied voltage, demonstrating the full tunability of the splitting ratio between output ports 1 and 2. With the measured input and output optical powers we also evaluated the SER and loss, as shown in Table 1. The loss and SER for the output port 1 is obtained by and for kV. Those for the output port 2 are obtained similarly for kV. Here, the loss at each output port includes the insertion loss () and the leakage to undesired output port (e.g., for the loss at the output port 1) caused by imperfect interference visibility (98%). For four input polarization states, the router has a loss of 2-4% and a SER of 20 dB. The slight difference in SER and loss for the output ports 1 and 2 are caused by the imbalanced transmission and reflection (49% versus 51%) of the beam splitters.
@@
Table 2 shows the performance of our router in comparison with previous works[16, 26, 27]. Our router has outperformed in loss, SER, and polarization process fidelity. Moreover, our router has possible improvements with feasible existing technologies. The absorption loss in the EOM ( 2%) can be mitigated at a longer wavelength; for example, at a telecommunication band, EOM with RTP crystals has less than 1% loss[6]. The 20 dB SER is currently limited by the imperfect (98%) interference visibility, which also introduces 1% loss at the desired output port. This may be improved by high-precision, high-stability optics and mechanical components. The EOM is compatible with a high-repetition-rate ( 1 MHz) driver designed for thermally compensated Pockels cells. The repetition rate can also be improved by using multiple pieces of our custom EOMs in the MZI. With the feasible improvement outlined above, the low-loss, high-speed routing of single-photon polarization qubits is possible for scaling up future photonic quantum information processing.

### 06 Apr 2024 v2

@@ 1 Introduction
In this work, we propose and experimentally demonstrate a low-loss, polarization-maintaining EO router compatible with single photons. Our scheme is based on the polarization-maintaining operation of all optical components including custom birefringence-compensated EO crystals. This enables the construction of a router with a Mach-Zehnder interferometer (MZI) configuration without polarization compensators. Our experiment with classical laser pulses demonstrates 2-4% loss, 20 dB switching extinction ratio (SER), and 99% process fidelity in the polarization degree of freedom. The measured routing rise time is as fast as a standard bulk optic EOM (2.9 ns). The EO Pockels effect employed in our experiment essentially does not produce noise photons, as demonstrated in previous quantum optics experiments.[6, 20, 21, 22, 23] Therefore, our router is expected to perform similarly for single-photon states having similar spectral and spatial modes to the classical light pulses used in our experiment. We expect that our demonstration is an important step toward advanced polarization-encoded photonic quantum information technologies that rely on photon-polarization qubits and their entanglement.
@@ 2 Methods
where denotes the wavelength of the incident light; and are the length and width of the RTP crystals; () and () are the intrinsic refractive indices and EO coefficient for crystallographic () direction, respectively; is the voltage applied to the crystal. For identical applied fields to the two crystals, i.e., , one can achieve the same phase shift for orthogonal polarization states. Thus, the static and EO birefringence of the two crystals is compensated for each other, enabling a polarization-maintaining phase shift according to the applied electric field. We note that our scheme is generally applicable to a pair of identical EO crystals that is often fabricated for off-the-shelf thermally compensated Pockels cells, and it applies even if the crystals have different widths and/or lengths, as long as the applied voltage to the crystals is calibrated accordingly. The other optical components of the router also preserve polarization states of light: in our squished configuration of the MZI, all optics components are operated with a nearly normal angle of incidence (AOI, ) to mitigate the polarization dependence (on S- and P-polarizations) in transmissivity, reflectivity, and phase shift[18, 19]. The interferometer arm without EOM is 16 mm longer than the other to compensate for the intrinsic group delay (relative to air) given by the EOM ( mm).
@@ 3 Results and Discussion
We constructed a polarization-maintaining router based on the scheme described above, using optical components with anti-reflection/high-reflection coatings optimized at the wavelength of nm. In this proof-of-concept experiment, we characterized our router via classical photodetection measurements of optical pulses from a diode laser at 785 nm with a bandwidth of 2 nm. The diode laser beam is coupled to a single-mode optical fiber and then collimated by an aspheric lens (with a focal length of 7.5 mm) so that an input optical pulse to the router has a Gaussian spatial mode. Our measurement predicts the router’s performance on single-photon states having similar bandwidth and spatial mode to the classical laser pulses since our router utilizes the EO effect which is essentially independent of input optical power as shown in Eq. (1, 2) and produces no noise photons, as has been demonstrated in many quantum optics experiments [6, 20, 21, 22, 23]. A set of a polarizer (POL), a half-wave plate (HWP), and a quarter-wave plate (QWP) is used to prepare an arbitrary input polarization state. The router is operated at a 10 kHz repetition rate, which is limited by our high-voltage EOM driver; we expect that a 1 MHz repetition rate is available with our EOM that is compatible with off-the-shelf high-voltage drivers designed for standard Pockels cells (used for polarization switching)[6, 16]. Our MZI is stable for several minutes which is limited by the environmental disturbance. Nevertheless, the stability can be improved by miniaturizing the setup and/or utilizing matured active phase stabilization techniques to be applied for advanced quantum optics experiments. In our experiment performed at the low repetition rate (10 kHz), the phase variation due to the temperature instability of the EOM is not observed. The temperature instability of the EOM that may occur at a higher repetition rate can be solved by active temperature stabilization techniques such as thermoelectric and water-cooling methods. A thermal compensation configuration can also be constructed by placing a polarization-maintaining EOM on each arm of the MZI and applying opposite voltages to them. This push-pull operation also reduces the necessary voltage for switching an output port by half.
@@
We measured optical power at the input port () and the output ports 1 and 2 ( and ) for different input polarization states and the applied voltage () to the EOM. Figure 2 shows the normalized output optical power ( and ) versus . The output power for horizontal (), diagonal (), and right-circular () polarized light changes according to the applied voltage, demonstrating the full tunability of the splitting ratio between output ports 1 and 2. We observed high interference visibility (98%) without spatial filtering. This indicates that the spatial mode of the input light is highly preserved in the router constructed by a minimum number of optics. With the measured input and output optical powers we also evaluated the SER and loss, as shown in Table 1. The loss and SER for the output port 1 are obtained by and for kV. Those for the output port 2 are obtained similarly for kV. Here, the loss at each output port includes the insertion loss () and the leakage to undesired output port (e.g., for the loss at the output port 1) caused by imperfect interference visibility (98%). For four input polarization states, the router has a loss of 2-4% and a SER of 20 dB. The slight difference in SER and loss for the output ports 1 and 2 are caused by the imbalanced transmission and reflection (49% versus 51%) of the beam splitters.
@@
Table 2 shows the performance of our router in comparison with previous works[16, 26, 27]. Our router has outperformed in loss, SER, and polarization process fidelity. Moreover, our router has possible improvements with feasible existing technologies. The absorption loss in the EOM ( 2%) can be mitigated at a longer wavelength; for example, at a telecommunication band, EOM with RTP crystals has less than 1% insertion loss[6]. The 20 dB SER is currently limited by the imperfect (98%) interference visibility, which also introduces 1% loss at the desired output port. This may be improved by high-precision, high-stability optics and mechanical components. Although we used optical pulses with a 2-nm bandwidth for characterizing the performance of our router, the acceptance bandwidth for < 1% degradation of the interference visibility is estimated to be 30 nm, with the group velocity dispersion (302 ) of the RTP crystals taken into account.[28]. The EOM is compatible with a high-repetition-rate ( 1 MHz) driver designed for thermally compensated Pockels cells. The repetition rate can also be improved by using multiple pieces of our custom EOMs in the MZI. With the feasible improvement outlined above, the low-loss, high-speed routing of single-photon polarization qubits is possible for scaling up future photonic quantum information processing.